A 0.125-kg hockey puck is sliding North across a frictionless ice surface at a constant speed of 39 m/s when a hockey stick exerts a force of 32 N [S25oE] on the puck for 0.15 s. Find the final velocity of the hockey puck
1
Expert's answer
2020-10-23T11:51:43-0400


The puch initialy has the momentum "\mathbf{p}_0" with the coordinates:


"\mathbf{p}_0 = (0,mv_0)"

where "m = 0.125kg" is the mass of the puck and "v_0 = 39m\/s" is its initial speed. The force exerted to the puck has the following coordinates:


"\mathbf{F} = F (\cos\theta, -\sin\theta )"

where "F = 32N" is its magnitude. According to the second Newton's law, this force exerts the following momentum change to the puck:


"\Delta\mathbf{ p} = \mathbf{F}\Delta t = (F \Delta t\cos\theta, -F \Delta t\sin\theta )"

where "\Delta t = 0.15s" is the time of interaction.

Thus, the final momentum will be:


"\mathbf{p} = \mathbf{p}_0+\Delta\mathbf{p}\\
\mathbf{p} = (0+F \Delta t\cos\theta,\space mv_0-F \Delta t\sin\theta) \approx(4.35, 2.85)kg\cdot m/s"

The final speed is:


"\mathbf{v} = \dfrac{\mathbf{p}}{m} = (34.8, 22.8)m/s"


Answer. "(34.8, 22.8)m\/s".